OpenStudy (yooooooo):

fins the derivative of: y=sin(1/x) y=ln(sqrert(x^2 +1) y=e^xcos2x y=x(lnx)^3

OpenStudy (anonymous):

Each of these require chain rule 1. sin is the outside function, 1/x is the inside function 2. ln is the outside, sqrt(x^2+1) is the first inside, (x^2+1) is the second inside 3. use product rule. e^x is the first function, cos(2x) is the second function. Don't forget to do chain rule with the 2x 4. again, product rule. x is your first function, (lnx)^3 is the second function. Again, chain rule with the (lnx)^3 If you need help with a specific one, let me know :)

OpenStudy (yooooooo):

so for 1 the answer would be cos 1/x^3?

OpenStudy (yooooooo):

and the second one would be x/square root of X^2+1) times x^2+1?

OpenStudy (anonymous):

I'll do the first one for you: d/dx sin(1/x) = cos(1/x) * d/dx (1/x) = cos(1/x) * -x^-2 = -cos(1/x)/x^2

OpenStudy (anonymous):

For the second: d/dx ln(sqrt(x^2+1) =1/sqrt(x^2+1)* d/dx sqrt(x^2+1) =1/sqrt(x^2+1)* (1/2)(x^2+1)^(-1/2)* d/dx (x^2+1) =1/sqrt(x^2+1)* (1/2)(x^2+1)^(-1/2)*(2x) which simplifies to =x/(x^2+1)

OpenStudy (anonymous):

check your answers in wolframalpha if you're not sure